$SPAD/src/input intlf.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool intlf.output
)set message test on
)set message auto off
)clear all
--S 1 of 2
exp(-x**2) * erf(x) / (erf(x)**3 - erf(x)**2 - erf(x) + 1)
--R
--R
--R                            2
--R                         - x
--R                 erf(x)%e
--R   (1)  ------------------------------
--R              3         2
--R        erf(x)  - erf(x)  - erf(x) + 1
--R                                                     Type: Expression Integer
--E 1
--S 2 of 2
integrate(%, x)
--R
--R
--R                     +---+    erf(x) - 1      +---+
--R        (erf(x) - 1)\|%pi log(----------) - 2\|%pi
--R                              erf(x) + 1
--R   (2)  -------------------------------------------
--R                        8erf(x) - 8
--R                                          Type: Union(Expression Integer,...)
--E 2
)spool
)lisp (bye)

———-

References^{[1] nothing}

4